OpenStudy (anonymous):

A line has a slope of and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct. A. -9x+4y=-56 B. 9x+4y=56 C. y=9 over 4x-14 D. y=9 over 4x+14 Will fan and give medal

OpenStudy (anonymous):

It has a slope of what?

OpenStudy (anonymous):

The question has a gap in it. Can you retype it?

OpenStudy (anonymous):

it has a slope of 9 over 4

OpenStudy (solomonzelman):

there are only 2 options with y-intercept of -14. re-write each answer choice in a slope intercept form, and tell me which options have a y-intercept -14.

OpenStudy (solomonzelman):

if you do user's work, I will report.

OpenStudy (solomonzelman):

y-intercept form is: y=mx+b

OpenStudy (solomonzelman):

re-write all te answer choice in that form. If you need help, I can guide you.

OpenStudy (anonymous):

No need to be rude Solomon

OpenStudy (anonymous):

we're all just trying to help out here

OpenStudy (solomonzelman):

I am just saying that if someone will do user's work I will report. because someone was almost about to. I am not being rude.

OpenStudy (anonymous):

Then next time you might want to ask nicely instead of threatening.

OpenStudy (anonymous):

Well, good luck to you two.

OpenStudy (solomonzelman):

I'll show you with the 1st equation. \(\large\color{black}{ -9x+4y=-56 }\) \(\large\color{black}{ -9x+4y \color{red}{+9x}=\color{red}{+9x}-56 }\) \(\large\color{black}{ 4y =9x-56 }\) \(\large\color{red}{ \frac{\color{black}{4y}}{4} =\frac{\color{black}{9x}}{4} -\frac{\color{black}{56}}{4} }\) \(\large\color{black}{ y =\frac{9}{4}x-14 }\)

OpenStudy (solomonzelman):

can you do this to all of your answer choices?

OpenStudy (anonymous):

I can try but i suck at math so......

OpenStudy (solomonzelman):

C and D are already in this form, you just need to do this thing to option B.

OpenStudy (solomonzelman):

option B says: 9x+4y=56 tell me what do you get after subtracting 9x from both sides?

OpenStudy (anonymous):

after subtracting the 9x from both sides the equation looked like this: 4y=-9x+56. does that look right???

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

so you have: 4y=-9x+56 now divide each thing in the equation by 4.

OpenStudy (solomonzelman):

okay answer 3 questions: what is 4y divided by 4? what is -9x divided by 4? what is 56 divided by 4?

OpenStudy (anonymous):

i got y=-9 over 4x+14

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

I got to go, tag someone else